OpenStudy (anonymous):

MEDAL! AND FAN! Solve the following equation, and check your solution. Be sure to show all work in order to receive full credit. Solve -17 + n/5 = 33.

OpenStudy (yoloshroom):

Get n/5 by itself, so add 17 to both sides. so the equation would then be n/5=33+17 Add it together so we get n/5=50 Then multiply both sides by 5 to remove the fraction from N, so n=50*5 and the answer so n=250

OpenStudy (anonymous):

thank you

OpenStudy (yoloshroom):

yup!